A farmer has available 400 yards of fencing and wishes to enclose a rectangular area. Express the area, A, of the rectangle as a function of the width, x, of the rectangle.
Answer by oberobic(2304)

x = width
y = length
.
p = perimeter
p = 2(x + y)
p = 400 yd
.
2(x+y) = 400
x+y = 200
y = 200-x
.
a = area
a = x*y
.
substitute for y
.
a = x(200-x)
a = 200x -x^2
.
done.